# Supplementary Information

### v1

@@ Discussion
We used QDs as a benchmarking sample, as it is widely used in photoblinking studies. In general, the blinking behavior of QDs is heavily influenced by factors such as their core composition (palstra2021python), shell thickness (efros2016origin), and the excitation power. These parameters influence the number and stability of accessible quasi-stable lifetime–intensity states. QD 605, used for the current study, has a CdSe core tuned to emit around 605 nm and capped with a thin ZnS shell with a thickness of <2<2 nm. Light absorption arises from excitonic transitions in the CdSe core, with multiple higher-energy transitions giving rise to significant broadening of the absorption toward the blue. The ZnS shell is transparent in this range and primarily serves to passivate any surface states and confine carriers. QDs with relatively thin shells generally follow Type-A blinking behavior (efros2016origin), which is driven by Auger recombination, i.e., recombination of an electron-hole pair that involves transfer of the energy to a different charge carrier instead of photon emission. This Type-A blinking is, therefore, associated with blinking from a lower-intensity, short-lifetime state (state 1) to a higher-intensity, longer-lifetime state (state 2), and hence the two states show a linear relation between the lifetime and intensity, in line with Figure 4C. The high density of the Auger-dominated state 1 arises because these datapoints occupy a narrow region in lifetime–intensity space, making the population appear tighter.
The results of the clustering protocol applied to the QD data indicated that two clusters were optimal (see Fig. 4E), and the weighted single-molecule average lifetime was found to be $7.3\,\pm 4.6\,\text{ns}$. Ref. (grecco2004ensemble) reports QD 605 lifetimes to be in the range of 102010-20 ns depending on excitation power and other experimental conditions, while the bulk (solution-phase) lifetime was measured as $6.3\pm 1.2\,\text{ns}$ in Ref. (bhuckory2016evaluating). Our experimental conditions were similar to those in the latter study, with QDs dispersed, no oxygen scavengers or additional additives such as glycerol used. Therefore, our measured average lifetime is in good agreement with the bulk value. The relatively high uncertainty in our lifetime arises primarily from the statistically small size of the dataset and the heterogeneity of the SM events. To fully establish the exact number of relevant states for this QD sample, one would have to systematically compare the switching dynamics and blinking behavior for different choices of the cluster number. Nevertheless, the two-cluster fitting is sufficient based on the switching dynamics (switching rates and frequencies) between these two clusters, which indicates a strong reversible blinking mechanism.
LHCII
LHCII is often modeled as a two-state system (gruber2018isolated, kruger2011fluorescence, kruger2011fluorescence2), though intermediate states have also been identified (kruger2011fluorescence, kruger2011fluorescence2). The LHCII sample analyzed in this study displayed a linear correlation between the intensity and lifetime, and the clustering protocol revealed a strongly emitting unquenched state, a quenched state, as well as an intermediate state (Fig. 5), consistent with previous studies (kruger2011fluorescence, kruger2011fluorescence2). The environmental conditions in our study closely matched those of Ref. (akhtar2015pigment), which reports ensemble measurements in solution without the use of oxygen scavengers, and found a bulk lifetime of 3.6ns3.6\ $\text${ns}, in good agreement with our measured long-lifetime component. In bulk studies, the average intensity per complex is orders of magnitude less than in SMS studies. Since blinking is strongly light-dependent (kruger2011fluorescence), it will contribute negligibly to the bulk lifetime, explaining why the bulk lifetime is expected to match the long-lifetime component in our study. The intermediate state (state $2$) is identified as the least stable of the three states (Table 2). If this sample were to be treated as a two-state model out of the box, it would distribute this intermediate data into both states and skew the data, indicating a perfect example of user bias. Instead, the clustering protocol, which helps with considerable noise reduction, reveals the underlying structure more accurately. This allows a user to potentially eliminate unwanted or weakly contributing clusters and analyze only the relevant clusters as if they were part of a robust, error-resistant measurement system.
PB
PB is a massive, multi-hundred-chromophoric system that is highly sensitive to excitation power and experimental conditions (gwizdala2016controlling). Since there was a large variation in the fluorescence intensities, the sample data had to be normalized to a global level prior to clustering. The excitation power was relatively low ($\sim 1.7$ nW), which led to a low fluorescence intensity switching rate (assefa2024apce), as evidenced from Table 3. The BIC analysis provides sufficient evidence that multiple clusters can be identified, with three clusters being the simplest model to describe the system. This, however, should be regarded as a minimal representation, since the lifetime–intensity distribution reveals much more complexity than only three populations, suggesting that the dynamics are far richer. In fact, eight subpopulations, the next candidate solution, are consistent with the compartmental model of van Stokkum and coworkers, where eight functional compartments were used to describe the excitation energy transfer dynamics in the core units of PB (see Fig. 9 in Ref. (van2018functional)). Importantly, their model was fully consistent with their time-resolved emission spectroscopy data. Information about higher-energy compartments can be assessed in particular for strong energy traps, in which case the emission is trap-limited and the fluorescence spectra are strongly blue-shifted (gwizdala2016controlling). Such trap-limited, pre-equilibration emission explains the clusters that show a non-linear relationship between the lifetime and intensity in Figure 6. However, this non-linearity is also partly attributed to the size heterogeneity that was not fully accounted for by our data normalization procedure. It, therefore, appears that eight clusters may be physically relevant for PB.
The longest measured lifetime component of the unquenched states is that of cluster 3, with an average weighted lifetime of $1.57\pm 0.19\,\text{ns}$, which matches well with the bulk lifetime of $\sim 1.6$ ns reported in Refs. (gwizdala2016controlling, squires2019single), where experimental conditions were similar. The expected similarity between the bulk lifetime and the long-lifetime component in our study is explained as for LHCII above, considering that the fluorescence intensity dynamics of PB are also strongly dependent on the excitation intensity (gwizdala2016controlling). The densest data of the long-lifetime component cluster, enclosed by the confidence ellipse, has a slightly longer lifetime ($1.57\pm 0.19\,\text{ns}$) than that of the cluster center ($\sim 1.5$ ns), and therefore marginally biases the average weighted lifetime upwards.
@@ Conclusions
While highly sensitive, SMS is prone to significant noise, which can easily introduce fitting errors, bias in the parameter estimates, and artificially inflated variability. This reduces the ability to identify true subpopulations, leading to higher uncertainty and weaker correlations. It is, therefore, crucial to perform careful and reliable noise reduction to preserve meaningful structure while minimizing the spurious effects of noise, without introducing new biases.
In this work, we highlighted the comparative advantages of using grouped levels over ungrouped resolved levels in data analysis, particularly for denoising and subpopulation identification. We also introduced an objective GMM clustering protocol based on a BIC scoring system to determine the optimal number of dense clustered regions in 2D SMS data. We specifically tested this clustering protocol on SM lifetime vs. intensity data, whereby we first benchmarked the capability of the protocol on simulated two- and four-state data, taking large intensity heterogeneity into account that may result from a case such as varying excitation probabilities of randomly oriented molecular dipoles. Thereafter, we tested the clustering protocol on Alexa 647 dye, QD 605, LHCII, and PB. We identified three data clusters for Alexa, LHCII, and PB, and two clusters for QD. We also deduced that the clustering protocol is useful in the identification (and thereby the removal) of clusters corresponding to misfits and nonoptimal sample preparation, and the assignment of minor clusters that are often obscured by noise. The clustering protocol typically reveals a non-monotonically decreasing plot for the BIC score per cluster, and therefore, several minima are expected. It is therefore recommended that a separate study be done to determine the effect of different identified optimal clusters on the switching dynamics.
We have considered only pairwise switching dynamics between statistically significant subpopulations, for which the number of switching pathways scales with $\frac{N(N-1)}{2}$. An alternative analysis could involve switches between any one state and all the other N1N-1 states together, which would decrease the number of switching rates to $2N$.
Although this study focused only on fluorescence lifetime–intensity correlations, the general nature of the clustering protocol makes it applicable to any two-dimensional parameter correlations. In addition, the protocol can readily be extended to higher dimensions.
Overall, it appears that grouping the resolved levels is necessary when the data is significantly broadened, and clustering further is even more necessary to identify physically relevant subpopulations. It is clear that the clustering protocol is capable of detecting clusters for erratic data (such as the Alexa data), statistically small data sets (such as with the QD data), as well as multistate data (LHCII and PB).
@@
The Supplementary Information contains additional supportive Figures.
@@ Acknowledgments
We wish to acknowledge Michal Gwizdala for the PB and LHCII isolations and the fruitful discussions. MACL was supported by the National Research Foundation (NRF), South Africa (PMDS22070633262). JLB was supported by the Vrije Universiteit Amsterdam-NRF Desmond Tutu programme. TPJK acknowledges funding from the NRF (grant nos. 137973 and 0403211945) and the Rental Pool Programme of the Council for Scientific and Industrial Research’s Photonics Centre, South Africa.
@@
All authors conceived the study. JLB and GTA performed the experiments, captured the data, and performed pre-clustering analyses. MACL developed the clustering software, performed the simulations, and analyzed the data. TPJK acquired funding, supervised research, and administered the project. MACL and TPJK wrote the manuscript. All authors read and approved the final manuscript.
@@
This Supplemental Information contains additional Figures as referenced in the main text.
@@
Simulated data additonal information
Figure S1: Outcome of the clustering protocol applied to the well-defined four-state data. A. Lifetime–intensity distribution of the data. B. Corresponding BIC score plot. C. Clustered four-state data with cluster centers indicated by black crosses. Ellipses are drawn around cluster centers, with a confidence of $0.9$.
Alexa Fluor additional information
Figure S2: Applying clustering protocol to the Alexa grouped data, using 5 clusters, to show the effect of overfitting. Ellipses are shown, with a confidence of $0.85$.
PB additional information
Figure S3: Outcome of applying the clustering protocol to the ungrouped data. A. BIC score plot for the ungrouped PB data. B. Corresponding clustering result using five clusters, with cluster centers indicated using black crosses. Representative ellipses are drawn to show which clusters correspond to which data, at a confidence of $0.85$.
Figure S4: Result of clustering the grouped PB data using eight clusters (cluster centers are shown as black crosses). Ellipses indicate the region of data points considered for each cluster, at a confidence of $0.85.$

### v2

@@ Discussion
We used QDs as a benchmarking sample, as they are widely used in photoblinking studies. In general, the blinking behavior of QDs is heavily influenced by factors such as their core composition (38), shell thickness (42), and the excitation power. These parameters influence the number and stability of accessible quasi-stable lifetime–intensity states. QD 605, used for the current study, has a CdSe core tuned to emit around 605 nm and capped with a thin ZnS shell with a thickness of <2<2 nm. Light absorption arises from excitonic transitions in the CdSe core, with multiple higher-energy transitions giving rise to significant broadening of the absorption toward the blue. The ZnS shell is transparent in this range and primarily serves to passivate any surface states and confine carriers. QDs with relatively thin shells generally follow Type-A blinking behavior (42), which is driven by Auger recombination, i.e., recombination of an electron-hole pair that involves transfer of the energy to a different charge carrier instead of photon emission. This Type-A blinking is, therefore, associated with blinking from a lower-intensity, short-lifetime state (state 1) to a higher-intensity, longer-lifetime state (state 2), and hence the two states show a linear relation between the lifetime and intensity, in line with Figure 6C. The high density of the Auger-dominated state 1 arises because these datapoints occupy a narrow region in lifetime–intensity space, making the population appear tighter. The results of the clustering protocol applied to the QD data indicated that two clusters were optimal (see Fig. 6E), and the weighted single-molecule average lifetime was found to be $7.3\,\pm 4.6\,\text{ns}$. Ref. (69) reports QD 605 lifetimes to be in the range of 102010-20 ns depending on excitation power and other experimental conditions, while the bulk (solution-phase) lifetime was measured as $6.3\pm 1.2\,\text{ns}$ in Ref. (70). Our experimental conditions were similar to those in the latter study, with QDs dispersed, no oxygen scavengers or additional additives such as glycerol used. Therefore, our measured average lifetime is in good agreement with the bulk value. The relatively high uncertainty in our lifetime arises primarily from the statistically small size of the dataset and the heterogeneity of the SM events. To fully establish the exact number of relevant states for this QD sample, one would have to systematically compare the switching dynamics and blinking behavior for different choices of the cluster number. Nevertheless, the two-cluster fitting is sufficient based on the switching dynamics (switching rates and frequencies) between these two clusters, which indicates a strong reversible blinking mechanism.
@@ Conclusions
While highly sensitive, SMS is prone to significant noise, which can easily introduce fitting errors, bias in the parameter estimates, and artificially inflated variability. This reduces the ability to identify true subpopulations, leading to higher uncertainty and weaker correlations. It is, therefore, crucial to perform careful and reliable noise reduction to preserve meaningful structure while minimizing the spurious effects of noise, without introducing new biases. Due to the specific structure of SMS data, our GMM and IC-based clustering protocol performs significantly better than K-means and DPGMM (a recommendable Bayesian method) in objectively identifying the statistically most likely number of clusters in noisy SMS datasets. We have benchmarked our protocol across a range of simulated and experimental datasets to validate its performance under diverse conditions.
For clean two- and four-state systems, the protocol reliably recovers the correct number of clusters and underlying state distributions, where cluster-quality metrics and agreement with ground truth are simultaneously optimized. Polarization and dipole-orientation effects introduce asymmetry and density shifts, producing apparent broadening without adding physical states; despite this, the protocol continues to identify statistically stable, well-supported partitions that reflect the dominant structure of the data. Recovery is robust for moderate state overlap (APGO<0.2<0.2), but at high overlap (APGO>0.4>0.4), individual states become statistically inseparable, leading the protocol to select fewer, broader clusters. Varying relative occupancies at fixed widths has a minimal effect on clustering unless extreme imbalances down-weight low-probability states, which are then merged while well-defined states remain accurately recovered. These results show that distribution width and geometric overlap are the primary determinants of clustering efficacy, with occupancy effects relevant only in extreme cases.
Across all simulation classes, including alternative blinking kinetics, clustering outcomes are largely insensitive to dwell-time distributions, underscoring the robustness of the protocol for diverse single-molecule datasets. In experimental datasets, where ground truth is inherently inaccessible, cluster selection must rely on statistically grounded criteria. Solutions exhibiting expected cluster tightness and appropriate data coverage provide the most defensible interpretation, even when alternative partitions exist. This emphasizes that statistical optimality is the appropriate objective for experimental single-molecule data. For Alexa 647, three clusters were identified in the ungrouped data, which reduced to two clusters after grouping, demonstrating that grouping improves denoising and suppresses fragmentation of noisy or sparsely populated states. For QD 605, two clusters were consistently recovered across simulations and experimental data, confirming stable detection of physically meaningful subpopulations.
Across all datasets, AIC and BIC generally exhibit consistent trends and minima, while ICL often follows a slightly different trajectory due to its stronger penalty for overlap; however, shared elbow points indicate that AIC or BIC alone may suffice for practical cluster selection. The protocol also facilitates identification and removal of misfit clusters, clusters arising from nonoptimal sample preparation, and minor clusters obscured by noise. Non-monotonic behavior of BIC and related IC scores further emphasizes that multiple local minima are expected; cluster selection should therefore be guided by a combination of IC trends and cluster-quality metrics rather than by a single global minimum. Follow-up studies exploring how different statistically defined clusters influence inferred switching dynamics would be valuable.
In this context, the goal of clustering is not the exhaustive reconstruction of every hypothetical underlying physical state, but the identification of statistically defensible, experimentally interpretable subpopulations supported by the available information. The proposed protocol is explicitly designed to operate within this constraint: it does not enforce a fixed cluster number, nor does it privilege formal global minima of information criteria when these correspond to over-partitioned or physically ambiguous solutions. Instead, it identifies a set of statistically plausible candidate solutions and evaluates them using independent measures of compactness and data coverage, thereby distinguishing genuine structure from artifacts of noise, overlap, or density imbalance. This distinction is particularly important for experimental SMS datasets, where the true number of states is fundamentally unknowable, and overfitting can easily masquerade as increased physical insight. By prioritizing statistical identifiability over nominal state count, the protocol provides a principled and minimally subjective framework for clustering noisy, heterogeneous single-molecule data. When the data support fine structure, it is recovered; when they do not, the method reports this limitation transparently rather than producing spurious resolution, especially when compared to other clustering methods. As such, the framework offers a robust foundation for downstream analyses of switching dynamics, state interconversion, and photophysical mechanisms across a broad range of single-molecule experiments.
@@
The Supplementary Information contains six supportive figures and one table.
@@ Acknowledgments
All authors conceived the study. BvH performed the experiments and captured the Alexa and QD data. JLB performed pre-clustering analyses. MACL developed the clustering software, performed the simulations, and analyzed the data. TPJK acquired funding, supervised research, and administered the project. MACL and TPJK wrote the manuscript. All authors read and approved the final manuscript.
@@
All authors conceived the study. BH performed the experiments and captured the data. JLB performed pre-clustering analyses. MACL developed the clustering software, performed the simulations, and analyzed the data. TPJK acquired funding, supervised research, and administered the project. MACL and TPJK wrote the manuscript. All authors read and approved the final manuscript.
@@
Figure S1: Clustering results for the four-state simulated lifetime–intensity dataset (300 particles, 10-min traces). A. Representative 40-ms binned intensity trace showing rapid photoblinking. B. Lifetime–intensity distribution of all traces, revealing four populations. C. IC score plots (AIC, BIC, and ICL) identify four clusters as optimal, with mean cluster tightness $\langle D^{2}\rangle\approx 2$ and FO $\sim 0.05$. D. Clustered distribution with centers at (500 cps, 0.8 ns), (2601 cps, 2.1 ns), (1205 cps, 2.4 ns), and (3100 cps, 1.1 ns) marked by crosses and enclosed by 0.95 confidence ellipses. ARI = 1, confirming accurate recovery of the simulated states.
Figure S2: Comparison of alternative clustering methods on a well-defined two-state simulated fluorescence lifetime–intensity dataset (10-min time, 300 particles). A. DPGMM with three prior components and a Dirichlet concentration of 0.1 assigns two clusters to the same high-intensity population, forming a low-occupancy “ghost” cluster; 0.95 confidence ellipses shown; computation time $\approx 2\times$ that of the GMM protocol. B. K-means elbow plot identifies two clusters as optimal; computation time $\approx 2\times$ that of GMM. C. K-means result for two clusters shows a visually reasonable separation of the two populations.
Figure S3: Clustering outcomes for simulated two-state lifetime–intensity datasets with different blinking models. A, C, E, G: Distributions for simulations with identical mean states but varied on-state dwell times—power-law (A), exponential (C), truncated power-law with $k=1.2$ (E), and truncated power-law with $k=3.5$ (G); off-state dwell times follow a power law in all cases. B, D, F, H: Corresponding clustered solutions with cluster centers (black crosses) and 0.95 confidence ellipses. Mean cluster tightness $\langle D^{2}\rangle\approx 2$, FO $\sim 0.05$, and ARI = 1, confirming accurate state recovery.
Figure S4: Comparison of clustering methods on a well-defined five-state simulated dataset (SD = 0.1 $\times$ mean; 10-min time, 300 particles). A. DPGMM with six prior components correctly recovers all five populations. B. DPGMM with eight priors assigns seven clusters, including two low-occupancy “ghost” clusters; 0.95 confidence ellipses shown, and Dirichlet weights set to 0.1. C. K-means elbow plot (red) with green gradient lines fitted to the steep and shallow regimes of the WCSS curve; the elbow is identified at their intersection, where the gradient transitions from rapid to slow decrease, yielding three or four plausible cluster numbers. D, E. K-means results for three (D) and four (E) clusters, showing poor recovery. F. K-means with five clusters explicitly specified, correctly recovering all populations. For both K-means and DPGMM, computation time was $\approx 2\times$ that of our GMM protocol.
Table S1: Means and standard deviations of the simulated intensity and lifetime parameters used to generate the ten-state, multi-state dataset composed of five well-defined states (1–5) and five diffuse, low-probability states (6–10).
State Mean intensity (cps) Intensity SD (cps) Mean Lifetime (ns) Lifetime SD (ns)
1 500 120 0.8 0.22
2 2600 200 2.1 0.30
3 1200 200 2.4 0.20
4 1500 200 1.4 0.20
5 2200 200 0.8 0.20
6 1900 600 1.0 0.40
7 1300 600 2.0 0.40
8 1000 600 0.6 0.40
9 3000 600 1.4 0.40
10 2000 600 2.2 0.40
Figure S5: Clustering results for a ten-state system (five well-defined and five diffuse states) under different blinking models. A–C: power-law on- and off-state dwell times; D–F: exponential on-state dwell times; G–I: truncated power-law on-state dwell times (exponent 1.2); J–L: truncated power-law on-state dwell times (exponent 3.5), with off-state dwell times following a power-law in all cases. Across all blinking models, three clusters are recovered at similar centers (approximately 510, 1170, and 2020 cps with lifetimes of 0.71, 2.3, and 1.45 ns), with mean cluster tightness $\langle D^{2}\rangle=1.80$–1.87, FO $\approx 0.03$, and ARI = 0.11–0.12. These results indicate that clustering performance in dense, overlap-dominated datasets is largely insensitive to the choice of blinking model.
Figure S6: Clustering of simulated FWHM versus peak-position data for multiple states. A. Scatter plot of the FWHM–peak position distribution. B. Ensemble-averaged intensity versus wavelength; C: Density plot of the FWHM–peak position distribution. D. IC score plots (0.95 confidence) identifying four clusters as optimal. E. Clustered result with centers (black crosses) and 0.95 confidence ellipses. ARI = 1, indicating perfect recovery of the simulated states.